### v1

@@ 1 Introduction
Large language models (LLMs) can follow natural language instructions, which has begun to reshape information retrieval (IR) [1]: an LLM receives a query, a set of candidate documents, and an instruction describing the ranking criterion; it then produces either a relevance judgment for a single document (pointwise [28, 29, 13, 17]) or a ranking over multiple documents (pairwise [19], setwise [30], listwise [15, 23, 24])[30].
@@
Prior work typically conducts a simple with-or-without comparison of the prompt with role-play phrasing versus a non–role-play baseline [22]. This leaves open an important question, especially as role-play becomes the standard prompting paradigm:
  • RQ0: How sensitive are LLM rankers to variations in role-play wording?
@@
To bridge this gap, we approach the problem from the perspective of mechanistic interpretability, which seeks to reverse-engineer the internal mechanisms of Transformer-based models. Specifically, we use activation patching [16, 25], a controlled intervention technique that replaces intermediate model activations with those from alternative prompts, allowing us to trace where and how information is represented and transmitted within the network. This allows us to address the following research questions:
@@ 3.1 Zero-shot LLM ranking with role-play
Zero-shot LLM-based ranking leverages the instruction-following capabilities of LLMs to assess query-document relevance via prompting. In this work, we consider both pointwise and pairwise zero-shot settings. In the pointwise setting, the model assesses one query-document pair at a time, which aligns naturally with relevance judgment tasks [28]. In the pairwise setting, each query is evaluated against two documents, a design that naturally encodes a relative relevance comparison and is therefore well-suited for ranking tasks [19]. To maintain consistency between pointwise and pairwise prompts, we unify the output format to “Yes” or “No” and structure each prompt according to the template illustrated in Figure 1 (a). For pairwise prompts, we control for positional bias by ensuring that Document A appears equally often as relevant and irrelevant across samples. Given the high computational cost of activation patching, we restrict each sample query to a single document (pointwise) or a single document pair (pairwise) and report results averaged over all samples.
@@ 3.2 Activation patching
In our setting, the behavioral contrast comes from role-play manipulation. A clean prompt instantiates a positive role (e.g., You are a capable search assistant…), while the corrupted counterpart instantiates a negative role (e.g., You are an unreliable search assistant…), as illustrated in Figure 1 (b). For both pointwise and pairwise ranking prompts, the expected label under the clean run for a relevant document is “Yes”, whereas the corrupted run for a relevant document is expected to produce “No”. By patching activations from the positive-role run into the negative-role run, we can test whether certain components are responsible for restoring the model’s ability to recognize relevance under the intended role. Formally, activation patching proceeds in three stages [27, 9]:
@@ 4.0.2 Dataset
We conduct most experiments on the MS MARCO passage ranking dataset [3] and use Natural Questions (NQ) [12] to evaluate generalization. Following the setup in Liu et al. [14], We sample 100 queries from the training set, each paired with one relevant document from the human-provided labels and one irrelevant document, both randomly drawn from the top 100 documents retrieved by BM25 [21]. For pointwise ranking, we use 50 samples with relevant documents and 50 with irrelevant ones. For pairwise ranking, we use 50 samples with Document A being relevant and 50 samples with Document A being irrelevant. Each Document A is paired with a counter-document (Document B) of the opposite relevance.
@@ 5.2 Information flow of role-play (RQ1)
Figure 3 shows that role-play information originates from role tokens in the lower layers, interacts moderately with instruction tokens in the middle layers, and gradually shifts toward the final token in higher layers. The mid-layer interaction is expected, as task-related information inevitably leaks into role specifications, making role and instruction difficult to fully disentangle. The shift to the final token is also consistent with prior findings [14, 16], since ranking scores are derived from the next-token logit at this position [14]. In contrast, patching document or query tokens yields negligible gains in both pointwise and pairwise methods. This implies that role-play signals embedded in the residual stream have limited interaction with query–document representations and are encoded largely independently of task-specific context.
We further examine the effect of prompt ordering on role-play encoding. As shown in Figure 4(a) , swapping the positions of document and query yields negligible changes, and patching these components consistently shows little effect. In contrast, Figure 4(b) shows that the relative position of role and instruction has only modest effects: when role tokens appear at the beginning, their influence remains mostly separable until mid-layers, whereas placing them immediately before the instruction induces slightly earlier role–instruction interaction and weak persistence into higher layers. In general, the flow of role-play information is largely insensitive to prompt component order, with only minor mid-layer interactions when the role precedes the instruction. We observe similar patterns for the Pairwise method.
@@ 5.3.2 Causal impact of attention heads across tasks
Specifically, we zero- or mean-ablate [26] the top-1 and top-10 attention heads (ranked by patching scores) for each prompt segment position – role, document, query, instruction, last token – and for mixed heads across these positions. Ablations are conducted on both pointwise relevance judgment and ranking tasks.
@@
We use the normalized probability of predicting the correct answer (correct score) to reflect model confidence, and this metric is preferred over accuracy, which reduces the task to a binary decision dependent on a threshold. Correct scores are expected to drop for both role types due to head ablation. Figure 6(a) shows that most scores for negative roles do drop (<0.393<0.393), while some scores for positive roles show a slight increase (>0.790>0.790). To further analyze this unexpected behavior, Figure 6(b) compares performance changes across different document types for positions of interest (role, instruction, and last token). We find that, for both role types, ablation effects are primarily driven by behavior on irrelevant documents. And the failure drop of irrelevant documents likely reflects the inherent difficulty in defining irrelevance consistently, as documents can be irrelevant in multiple ways.
@@
We further evaluate whether heads identified on MSMARCO generalize to TREC DL 2019. Correct scores are used to rerank the top-10 BM25 candidates, with NDCG@10 expected to decrease for positive roles and increase for negative roles. Figure 6 shows that the patterns are consistent with the relevance judgment task but more stable, likely because ranking evaluates multiple documents per query and reduces pointwise noise. Overall, the results indicate that the identified heads generalize well from single-document relevance judgments to the multiple-document ranking task.
@@ 5.5 Role-play encoding across models and datasets(RQ4)
Generalization across different models
We extend activation patching to other instruction-tuned LLMs: Mistral-7B-Instruct-v0.3 [10] and Qwen2.5-7B-Instruct [20], which are comparable in scale and architecture to LLaMa. As shown in Figure 9, both models reproduce the core information flow pattern, with role signals gradually propagating to the final token. At the same time, model-specific differences emerge: Mistral exhibits earlier and more sustained role–instruction interaction, while Qwen shows a delayed transition, beginning only around layer 20. Under the Pairwise setting, both models also display slightly stronger but still limited interaction between role and query–document tokens, likely because comparative judgments naturally leverage more contextual signals. Overall, the results suggest that while role encoding is stable across instruction-tuned LLMs, models differ in the timing and extent of role–instruction interaction, potentially reflecting architectural or training differences.
Generalization across different datasets
We also conduct our analyses on the NQ dataset using LLaMa-3.1-8B-Instruct, shown in the last column of Figure 9. The results reveal patterns highly consistent with those observed on MS MARCO: role signals dominate, interact modestly with instruction tokens, and show minimal entanglement with query–document content. This consistency suggests that the mechanism underlying role encoding is largely stable across datasets, pointing to an intrinsic property of the model’s representation rather than being tightly coupled to any specific data distribution.
@@ 6 Conclusions
In this paper, we have investigated the effect of role-play in LLM rankers, both through empirical evaluation of positive and negative role-play prompts and through the lens of mechanistic interpretability. We found that the formulation and wording of the role-play prompt has a large effect on the zero-shot ranking quality of the LLM. Model activations indicate that role-play signals interact with the instruction part of the prompt and hardly with the query and document representations. Our results generalize across multiple instruction-tuned 8B models and between pointwise and pairwise approaches. Our findings show that it is important that role-play prompts for LLM rankers are carefully constructed, extending findings from previous work that the effect of prompting can even be larger than the ranking paradigm (pointwise or pairwise). Our work can be further extended to more models, more datasets, listwise and setwise rankers, and specific domains that may require specific roles.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated a remarkable ability to follow natural language instructions, offering new paradigms for information retrieval (IR) [1]: an LLM receives a query, a set of candidate documents, and an instruction describing the ranking criterion; it then produces either a relevance judgment for a single document (pointwise [14, 18, 29, 30]) or a ranking over multiple documents (pairwise [20], setwise [31], listwise [16, 24, 25])[31].
@@
Prior work typically conducts a simple with-or-without comparison of the prompt with role-play phrasing versus a non-role-play baseline [23]. This leaves open an important question, especially given the increasing prevalence of role-play as a common prompting strategy [12]:
  • RQ0: How sensitive are LLM rankers to variations in role-play wordings?
@@
To bridge this gap, we approach the problem from the perspective of mechanistic interpretability, which seeks to reverse-engineer the internal weights and representations of neural models into human-understandable algorithms. Specifically, we use activation patching [17, 26], a controlled intervention technique that replaces intermediate model activations with those from alternative prompts, allowing us to trace where and how information is represented and transmitted within the network. This allows us to address the following research questions:
@@ 3.1 Zero-shot LLM ranking with role-play
Zero-shot LLM-based ranking leverages the instruction-following capabilities of LLMs to assess query-document relevance via prompting. In this work, we consider both pointwise and pairwise zero-shot settings. In the pointwise setting, the model assesses one query-document pair at a time, which aligns naturally with relevance judgment tasks [29]. In the pairwise setting, each query is evaluated against two documents, a design that naturally encodes a relative relevance comparison and is therefore well-suited for ranking tasks [20]. We exclude listwise and setwise approaches primarily because their multi-document context introduces intricate cross-document attention dependencies that would complicate the causal tracing of internal mechanisms. Furthermore, such settings are less amenable to a constrained and unified output space. To maintain consistency between pointwise and pairwise prompts, we unify the output format to “Yes” or “No” and structure each prompt according to the template illustrated in Figure 1a. For pairwise prompts, we control for positional bias by ensuring that Document A appears equally often as relevant and irrelevant across samples. Given the high computational cost of activation patching, we restrict each sample query to a single document (pointwise) or a single document pair (pairwise) and report results averaged over all samples.
@@ 3.2 Activation patching
In our setting, the behavioral contrast comes from role-play manipulation. A clean prompt instantiates a positive role (e.g., You are a capable search assistant…), while the corrupted counterpart instantiates a negative role (e.g., You are an unreliable search assistant…), as illustrated in Figure 1b. For both pointwise and pairwise ranking prompts, the expected label under the clean run for a relevant document is “Yes”, whereas the corrupted run for a relevant document is expected to produce “No”. By patching activations from the positive-role run into the negative-role run, we can test whether certain components are responsible for restoring the model’s ability to recognize relevance under the intended role. Formally, activation patching proceeds in three stages [9, 28]:
@@ 4.0.2 Dataset
We first assess the sensitivity of zero-shot LLM rankers to role-play wordings using the TREC-DL datasets [5]. For activation patching, we primarily utilize MS MARCO [3] for relevance judgment tasks, while Natural Questions (NQ) [13] is used to evaluate generalization. Additionally, ranking-specific patching tasks are conducted on TREC-DL. Following the patching setup in Liu et al. [15], we sample 100 queries from the training set, each paired with one relevant document from the human-provided labels and one irrelevant document, both randomly drawn from the top 100 documents retrieved by BM25 [22]. For pointwise ranking, we use 50 samples with relevant documents and 50 with irrelevant ones. For pairwise ranking, we use 50 samples with Document A being relevant and 50 samples with Document A being irrelevant. Each Document A is paired with a counter-document (Document B) of the opposite relevance.
@@ 5.2 Information flow of role-play (RQ1)
Figure 3 shows that role-play information originates from role tokens in the lower layers, interacts moderately with instruction tokens in the middle layers, and gradually shifts toward the final token in higher layers. The mid-layer interaction is expected, as task-related information inevitably leaks into role specifications, making role and instruction difficult to fully disentangle. The shift to the final token is also consistent with prior findings [15, 17], since ranking scores are derived from the next-token logit at this position [15]. In contrast, patching document or query tokens yields negligible gains in both pointwise and pairwise methods. This implies that role-play signals embedded in the residual stream have limited interaction with query-document representations and are encoded largely independently of task-specific context.
We further examine the effect of prompt ordering on role-play encoding. As shown in Figure 4(a) , swapping the positions of document and query yields negligible changes, and patching these components consistently shows little effect. In contrast, Figure 4(b) shows that the relative position of role and instruction has only modest effects: when role tokens appear at the beginning, their influence remains mostly separable until mid-layers, whereas placing them immediately before the instruction induces slightly earlier role-instruction interaction and weak persistence into higher layers. In general, the flow of role-play information is largely insensitive to prompt component order, with only minor mid-layer interactions when the role precedes the instruction. We observe similar patterns for the Pairwise method.
@@ 5.3.2 Causal impact of attention heads across tasks
Specifically, we zero- or mean-ablate [27] the top-1 and top-10 attention heads (ranked by patching scores) for each prompt segment position – role, document, query, instruction, last token – and for mixed heads across these positions. We report the top-10 results as a representative summary. Ablations are conducted on both pointwise relevance judgment and ranking tasks.
@@
This task evaluates the model’s ability to independently assess the relevance of individual query-document pairs. We use the normalized probability of predicting the correct answer (correct score) to reflect model confidence, and this metric is preferred over accuracy, which reduces the task to a binary decision dependent on a threshold. Correct scores are expected to drop for both role types due to head ablation. Figure 6(a) shows that most scores for negative roles do drop (<0.393<0.393), while some scores for positive roles show a slight increase (>0.790>0.790). To further analyze this unexpected behavior, Figure 6(b) compares performance changes across different document types for positions of interest (role, instruction, and last token). We find that, for both role types, ablation effects are primarily driven by behavior on irrelevant documents. And the failure drop of irrelevant documents likely reflects the inherent difficulty in defining irrelevance consistently, as documents can be irrelevant in multiple ways.
@@
In contrast, the ranking task involves the relative comparison of individual relevance judgments to determine the final order of multiple documents for a given query. We further evaluate whether heads identified on MSMARCO generalize to TREC DL 2019. Correct scores are used to rerank the top-10 BM25 candidates, with NDCG@10 expected to decrease for positive roles and increase for negative roles. Figure 7 shows that the patterns are consistent with the relevance judgment task but more stable, likely because ranking evaluates multiple documents per query and reduces pointwise noise. Overall, the results indicate that the identified heads generalize well from single-document relevance judgments to the multiple-document ranking task.
@@ 5.5 Role-play encoding across models and datasets (RQ4)
5.5.1 Generalization across different models
We extend activation patching to other instruction-tuned LLMs that share similar scales and architectures: Mistral-7B-Instruct-v0.3 [10] and Qwen2.5-7B-Instruct [21]. As shown in Figure 9, both models reproduce the core information flow pattern, with role signals gradually propagating to the final token. At the same time, model-specific differences emerge: Mistral exhibits earlier and more sustained role-instruction interaction; while Qwen shows a delayed transition to the last token, beginning only around layer 20. Under the pairwise setting, both models also display slightly stronger but still limited interaction between role and query-document tokens, likely because comparative judgments naturally leverage more contextual signals. Overall, the results suggest that while role encoding is stable across instruction-tuned LLMs, models differ in the timing and extent of role-instruction interaction, potentially reflecting architectural or training differences.
5.5.2 Generalization across different datasets
We also conduct our analyses on the NQ dataset using LLaMa-3.1-8B-Instruct, shown in the last column of Figure 9. The results reveal patterns highly consistent with those observed on MS MARCO: role signals dominate, interact modestly with instruction tokens, and show minimal entanglement with query-document content. This consistency suggests that the mechanism underlying role encoding is largely stable across datasets, pointing to an intrinsic property of the model’s representation rather than being tightly coupled to any specific data distribution.
6 Discussion
While our analysis provides insight into role-play mechanisms, the choice of role descriptors was influenced by activation patching requirements. The need for precise position-wise alignment led us to focus on single-token wordings. Since single-token ranking-related descriptors are scarce, we also included more general or unrelated negative terms (such as ‘awful’ and ‘sadly’). The fact that such descriptors can still degrade ranking quality indicates that role-play effects need not be semantically tied to ranking expertise, motivating broader mechanistic analysis to prevent unintended or adversarial triggers of ranking degradation. Our mechanistic analysis provides a roadmap for model steering. Patching reveals that role-play signals are processed early and propagated to the final token in later stages, enabling targeted layer and position interventions, while their weak interaction with query-document representations suggests that role-play can be steered independently without disrupting core relevance, supporting guardrails against negative role leakage and role-based ranking improvements.
@@ 7 Conclusions
In this paper, we have investigated the effect of role-play in LLM rankers, both through empirical evaluation of positive and negative role-play prompts and through the lens of mechanistic interpretability. We found that the formulation and wording of the role-play prompt has a large effect on the zero-shot ranking quality of the LLM. Model activations indicate that role-play signals interact with the instruction part of the prompt and hardly with the query and document representations. Our results generalize across multiple instruction-tuned 8B models and between pointwise and pairwise approaches. Our findings show that it is important that role-play prompts for LLM rankers are carefully constructed, extending findings from previous work that the effect of prompting can even be larger than the ranking paradigm (pointwise or pairwise). Our work can be further extended to more models, more datasets, and specific domains that may require specific roles.
{credits}
7.0.1 Acknowledgements
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.